OpenStudy (anonymous):

what conic section represents this equation: x^2+y^2-6x+8y-24=0

OpenStudy (anonymous):

İF YOU WANT I CAN SHOW

OpenStudy (anonymous):

x^2+y^2-6x+8y=24 AND x^2-6x+(9)+y^2+8y+(16)=24+25

OpenStudy (anonymous):

(x-3)^2+ (y+4)^2=49

OpenStudy (anonymous):

so it a circle center ( 3,-4) r= 7

OpenStudy (anonymous):



OpenStudy (anonymous):

@chillingsy